use long division and check (x^3-64)divided by(x-4)
Answer by ewatrrr(24785)

 
Hi,
x-4 |x^3 - 64 is x^2 +4x +16

and...(x-4)(x^2+4x+16) = x(x^2+4x+16) - 4(x^2+4x+16) = x^3 - 64